4. The areas of the bases of a frustum of a regular pyramid is 250 m2 and 640 m2 respectively. Find the altitude of the pyramid if the altitude of its frustum is 130 m?
1
Expert's answer
2020-12-10T09:59:51-0500

"A .S.F=\frac{640}{250}"

This implies that "A.S.F=\frac{64}{25}"

"L.S.F= \sqrt{\frac{64}{25}}"

Therefore, "L.S.F=\frac{8}{5}"

"\frac{X+130}{X}=\frac{8}{5}"

Solving for X,

"5X+650=8X"

"3X=650"

"X=216.67m"

Therefore, the altitude of the pyramid is: "X+130=346.67m"